keke2003:

If m(x)=x+5/x-1 and n(x)=x-3, which function has the same domain as (mon)(x)? o means composition in (mon)

jhonyy9:

do you know this rule for composition of functions ? (f ° g)(x) = f(g(x))

keke2003:

f(g(x))=f(x^2)=x^2-7

jhonyy9:

so (m°n)(x) = m(n(x)) what mean that in equation of m(x) in place of x substitute the equation of n(x) please post your work